The 100-m record is currently held by Usain bolt with a time of 9.8 seconds. What is Bolt's average speed over the 100-m? A bell of natural frequency 352-Hz is being rung at the finish line as bolt approaches it. What frequency does Bolt hear? What is the bell's natural (rest) wavelength? If the bell rung at 80 dB, what is the intensity level?
Answer #1

Average speed of bolt = 100/9.8 = 10.20 m/s

The apparent frequency heard by bolt, ν’ = ν(v – vL/v – vS) = 352(344 + 10.20/344 – 0) = 362.437 Hz

The natural wavelength of bell, λ = v/ν = 344/352 = 0.977 m

As the sound level, β = 10 dB log (I/I0)

80 dB = 10 dB log (I/10-12)

8 = log (I/10-12)

I = 10-4(W/m2)